Two teachers have classes of similar sizes. After the final exams, the mean of the grades in each class is 73%. However, one class has a standard deviation of 4% while the other is 8%. In which class would a mark of 90% be more meaningful?
1
Expert's answer
2020-05-24T20:20:38-0400

the distribution in both classes can be taken as Normal distribution.

A1 first class, A2 second class.

A1=N(73,4), A2=N(73,8)

P1={A1>90}= "1-\Phi((90-73)/4)=1-\Phi(4.25)=0"

"\Phi(3,5)=1"

P2={A2>90}="1-\Phi((90-73)/8)=1-\Phi(2.125)=1-0.9834=0.0166"

this means that in the second grade, there was a probability of about 2% of students with a score of more than 90%, when in the first there are no such students at all

in the second grade a score of 90% is more likely

the reference to a calculation formula and a data table

(https://en.wikipedia.org/wiki/Normal_distribution,https://www.normaltable.com)